Gas Density and Molar Mass:
Part A
Calculate the density of oxygen, O2, under each of the following conditions:
STP
1.00 atm and 30.0 ∘C
Express your answers numerically in grams per liter. Enter the density at STP first and separate your answers by a comma.

Part B
To identify a diatomic gas (X2), a researcher carried out the following experiment: She weighed an empty 4.5-L bulb, then filled it with the gas at 1.70 atm and 26.0 ∘C and weighed it again. The difference in mass was 8.7 g . Identify the gas.
Express your answer as a chemical formula.
1
Expert's answer
2017-03-21T10:02:06-0400
Part A

pV=mRT/M; m/V-pM/RT; p=101325Pa;T=303ᵒK

m/V=101325Pa*0.032kg/Mol/8.314J/Mol*ᵒK*303ᵒK=1.29 kg/M³, 1.29g/L


Part B

V=0.0045m³;p=172252.5Pa;T=299ᵒK;m=0.0087kg

M=mRT/pV=0.0087kg*8.314JMolᵒK*299K/172252.5Pa*0.0045m³=0.0279kg/Mol=28g/Mol

Unknown metal can be N₂;CO;C₂NH